# Optimal Bias-Correction and Valid Inference in High-Dimensional Ridge Regression: A Closed-Form Solution

### v1

@@ 3   Monte Carlo Simulations
In this section, we conduct Monte-Carlo experiments to illustrate the proposed procedure. We consider two scenarios where $p<n$ and $p>n$ and study the effect of the bias-correction procedure and the approximation of the asymptotic normalities established in Section 2.5.
Example 1. In this example, we consider the data-generating process in (2) for different settings of $(p,n)$ with $p<n$. For each configuration of $(p,n)$, we set the seed number in R software as (1234) and first generate an $n\times p$ matrix ${\mathbf{H}}$, where its elements are drawn independently from a Uniform distribution $U(-2,2)$. We then perform a singular-value decomposition on ${\mathbf{H}}$ and obtain its left and right singular vectors ${\mathbf{M}}\in R^{n\times p}$ and ${\mathbf{N}}\in R^{p\times p}$, then we let ${\mathbf{X}}={\mathbf{M}}{\mathbf{N}}^{\prime}$. The first $p/2$ elements of $\boldsymbol{\beta}$ are generated from $U(-2,-1)$ independently, and the remaining $p/2$ elements are from $U(1,2)$. In each replication, the noise $\varepsilon$ is generated from multivariate normal distribution $N({\bf 0},{\mathbf{I}}_{n})$. We consider the scenarios where $(p,n)=(50,100)$, $(50,400)$, $(100,200)$, and $(100,500)$. The choices of $\lambda$ are $\lambda=0.05n$, $0.1n$, $0.3n$, and $0.5n$. 1000 replications are used in each setting throughout the experiments.
In Table 1, we report the estimation errors of the ridge estimators and the de-biased ones. For each configuration of $(p,n,\lambda)$, we define the empirical mean squared errors (MSEs) as
@@
where $\widehat{\boldsymbol{\beta}}^{(j)}(\lambda)$ and $\widehat{\boldsymbol{\beta}}_{c,k}^{(j)}(\lambda)$ are the ridge estimators and the de-biased one with $k$ iterations in the $j$-th replication, respectively. The standard errors $\widehat{\sigma}_{0}$ and $\widehat{\sigma}_{100}$ are estimated via
$\widehat{\sigma}_{0}=\sqrt{\frac{1}{1000}\sum_{j=1}^{1000}\frac{1}{n}\|{\mathbf{y}}-{\mathbf{X}}\widehat{\boldsymbol{\beta}}^{(j)}(\lambda)\|_{2}^{2}}\,\,\text{and}\,\,\widehat{\sigma}_{100}=\sqrt{\frac{1}{1000}\sum_{j=1}^{1000}\frac{1}{n}\|{\mathbf{y}}-{\mathbf{X}}\widehat{\boldsymbol{\beta}}_{c,100}^{(j)}(\lambda)\|_{2}^{2}}.$ (43)
In other words, we evaluate the estimations of the standard errors by the ridge estimators and the de-biased ones.
For each setting in Table 1, we see that the MSE first decreases as the number of iterations increases, and then increases as we use more iterations to correction the bias terms. This is understandable because it is in line with the bias-variance trade-off in the machine learning literature. See, for example, Hastie et al. (2009). Specifically, we plot the MSEs for $(p,n)=(50,100)$ and $(100,200)$ with $\lambda=0.05n$ in Figure 1, where we can clearly see that the MSEs can achieve a minimum point as we increase the number of iterations, or equivalently, the model complexity, and the MSE will further increase as we try to completely correct the biases. Finally, the MSE becomes stable, which is consistent with our asymptotic theory. Moreover, the standard error estimated using the debiased ridge estimator is closer to the true one (unity), showing the effectiveness of our bias-correction procedure.
@@
Furthermore, we study the consistency of the de-biased estimators using our proposed method. For simplicity, we investigate the averaging estimation errors (AEEs) of the ridge estimators and the de-biased ones, which are defined, respectively, as
@@
where the ones in the $\ell_{2}$-norm are the empirical versions of the biases. The AEEs are reported in Table 2. From Table 2, we can see a decreasing pattern for each $(p,n)$ as the number of iterations increases, implying that our de-biased estimators are consistent to the true parameters for sufficiently large $k$.
Table 2: Empirical averaging estimation errors (AEEs) when $p<n$ in Example 1, where the AEEs are defined in (44) and (45) for ${\mathbf{b}}_{\lambda}$ and ${\mathbf{b}}_{\lambda,k}$, respectively. The number of iterations $k=1,5,10,20,50$, and $100$. 1000 replications are used in the experiments.
@@
Finally, we investigate the performance of the inference approach based on the de-biased estimators in Theorem 6 for $p<n$. For simplicity, we only consider the case when $(p,n)=(100,200)$ and $\lambda=0.3n$, and we can produce similar results for other settings. Let ${\mathbf{e}}_{i}$ be the $i$-th standard unit vector of $R^{p}$,
@@
where ${\bf 0}_{s}$ is an $s$-dimensional vector of zeros. Figure 2 presents the histograms of $\sqrt{n}{\mathbf{e}}_{1}^{\prime}(\widehat{\boldsymbol{\beta}}(\lambda)-\boldsymbol{\beta})$, $\sqrt{n}{\mathbf{e}}_{2}^{\prime}(\widehat{\boldsymbol{\beta}}(\lambda)-\boldsymbol{\beta})$, $\sqrt{n}\boldsymbol{\theta}_{1}^{\prime}(\widehat{\boldsymbol{\beta}}(\lambda)-\boldsymbol{\beta})$, and $\sqrt{n}\boldsymbol{\theta}_{2}^{\prime}(\widehat{\boldsymbol{\beta}}(\lambda)-\boldsymbol{\beta})$ obtained from 1000 experiments. From Figure 2, we see that the biases of the empirical means in all the histogram plots are significantly large and diverges from zero, implying that the traditional ridge estimators are not appropriate to make statistical inferences without the information of the bias terms. In contrast, we plot the empirical histograms of the de-biased counterparts in Figure 3 under the same setting. From Figure 3, we see that the finite sample performance is quite satisfactory, and the bias effect of the de-biased estimators is very small. In addition, the curve of the normal distribution in Theorem 6(i) is added to the corresponding histograms, where the standard error term $\sigma$ is estimated from the data as that in Remark 5(iv) and Table 1. From these curves, we can further confirm that our inference method is valid for the de-biased ridge estimators, indicating that valid inferences based on the ridge estimators with our bias-correction method can be made in practice.
@@
Figure 3: Empirical histograms of (a) $\sqrt{n}{\mathbf{e}}_{1}^{\prime}(\widehat{\boldsymbol{\beta}}_{c,k}(\lambda)-\boldsymbol{\beta})$; (b) $\sqrt{n}{\mathbf{e}}_{2}^{\prime}(\widehat{\boldsymbol{\beta}}_{c,k}(\lambda)-\boldsymbol{\beta})$; (c) $\sqrt{n}\boldsymbol{\theta}_{1}^{\prime}(\widehat{\boldsymbol{\beta}}_{c,k}(\lambda)-\boldsymbol{\beta})$; and (d) $\sqrt{n}\boldsymbol{\theta}_{2}^{\prime}(\widehat{\boldsymbol{\beta}}_{c,k}(\lambda)-\boldsymbol{\beta})$ in Example 1. We choose $k=120$ for the de-biased estimators with $(p,n)=(100,200)$ and $\lambda=0.3n$. A density curve is plotted based on the limiting distribution outlined in Theorem 6(i). 1000 replications are used in the experiments.
@@
Example 2. In this example, we consider the scenario where $p>n$. The seed number used in this example is the same as that in Example 1. Each row of the design matrix ${\mathbf{X}}$ is generated independently from multivariate standard normal distribution $N({\bf 0},{\mathbf{I}}_{p})$. For each configuration of $(p,n)$, the first $5$ elements in $\boldsymbol{\beta}$ are generated from $U(-5,-2)$, the $6$th to the $10$th element are from $U(2,5)$, and the remaining ones are zeros. In other words, we consider the sparse vector $\boldsymbol{\beta}$ with $p^{*}=10$ in (18). The dimensions and the sample sizes are $(p,n)=(150,120),(150,140),(220,180)$, and $(220,200)$, and we consider $\lambda=0.1n$, $0.3n$, and $0.8n$ in each replication for any given $(p,n)$. For simplicity, we choose $n^{*}=40$ in the ridge screening of (20) with $k=100$ iterations. $1000$ replications are used for each configuration of $(p,n)$ throughout the experiments.
@@
In Table 3, we report the empirical MSEs of the ridge estimators and the de-biased ones before applying the ridge-screening method, and those after the ridge-screening with $k=100$ iterations in the first stage of bias-corrections. Specifically, the upper panel in Table 3 for each $\lambda$ presents the MSEs of the ridge estimators and the de-biased ones using 100 iterations without applying the ridge-screening approach. We see that the MSEs of the de-biased estimators become quite stable as we increase the number of iterations, which is consistent with our findings that there is a term that cannot be corrected from the data. In addition, the estimated standard errors are not close to unity, which is the true one in the experiments. In the lower panel of Table 3 for each $\lambda$, we apply the ridge-screening approach with $k=100$ and sort out the largest $n^{*}=40$ components of $\widehat{\boldsymbol{\beta}}_{c,100}(\lambda)$, and then obtain the restricted ridge estimators. We further apply the bias-correction approach and the de-biased estimators are obtained in another $100$ iterations. From the lower panels of Table 3 for each $\lambda$, we see that the bias of the de-biased estimators decreases sharply to a relatively stable value for each configuration of $(p,n)$, which is understandable since the selected model consists of more parameters than those in the true one. In addition, the estimation errors are significantly smaller than those without the ridge-screening approach, and the estimated standard errors after bias correction are closer to the true one than the methods without using the ridge-screening and the bias-correction approaches. We also note that the estimated standard errors are still larger than the true one (unity) when $\lambda=0.8n$ after the ridge-screening, but they are much closer to the true ones than all the ones produced by the original ridge estimators. In addition, this can be optimized by choosing more appropriate tuning parameter $\lambda$ (e.g. $\lambda=0.1n$ or $0.3n$) such that the estimated standard errors are close to one. Furthermore, we also note that there is also a bias-variance trade-off in the MSEs of the de-biased estimators with or without the ridge-screening. Figure 4 plots the the empirical MSEs of the de-biased ridge estimators for different number of iterations for $\lambda=0.5n$ and $(p,n)=(150,120)$, where (a) plots the MSEs of the de-baised ridge estimators before applying the ridge-screening method, and (b) provides the MSEs of the de-biased ridge estimators after the ridge-screening. We can clearly see that there is an optimal number of iteration that minimizes the MSE, which is in line with our asymptotic results in Theorem 7.
@@
We also study the performance of the proposed ridge-screening method in Table 4, where the empirical probability (EP) is calculated by
@@
where $\mathcal{M}_{k}^{i}(\lambda^{*})$ is the recovery in the $i$-th experiment and $|\mathcal{M}_{0}\cap\mathcal{M}_{k}^{i}(\lambda^{*})|$ is the cardinality of the set $\mathcal{M}_{0}\cap\mathcal{M}_{k}^{i}(\lambda^{*})$. We can see from Table 4 that our method provides satisfactory performance in variable selections [^3] The comparison results with other variable selection approaches, such as the LASSO, are not provided in this experiment due to the empirical probability of correct recoveries by the proposed RS method being $100\%$ in all settings of Table 4..
@@
Figure 4: The bias-variance trade-off reflected by the empirical MSEs of the de-biased ridge estimators for different number of iterations in Example 2, where we consider $\lambda=0.5n$ for $(p,n)=(150,120)$, where (a) plots the MSEs of the de-baised ridge estimators before applying the ridge-screening method, and (b) provides the MSEs of the de-biased ridge estimators after the ridge-screening. 1000 replications are used in the experiments.
Table 4: Empirical probability (EP) of correct recoveries of the true models for $\lambda^{*}=0.1n$, $0.3n$, and $0.8n$, where we conduct 100 iterations in the bias-correction procedure and set $n^{*}=40$ in the ridge-screening. 1000 replications are used in the experiments.
$\lambda^{*}=0.1n,k=100$
$(p,n)$
EP (150,200) (150,140) (220,180) (220,200)
$\lambda^{*}=0.3n,k=100$
$(p,n)$
EP (150,200) (150,140) (220,180) (220,200)
$\lambda^{*}=0.8n,k=100$
$(p,n)$
EP (150,200) (150,140) (220,180) (220,200)
Similar to the experiments in Example 1, we present the empirical averaging estimation errors (AEEs) in Table 5. The settings of $(p,n)$, $\lambda^{*}$, $\lambda$, $k$, and $l$ are the same as those in Table 3. From Table 5, we see that the averaging estimation errors before ridge-screening are also quite stable as we increase the number of iterations which is similar to the findings in Table 3. After we apply the ridge-screening approach to the de-biased estimators in the upper panel of Table 5 with $k=100$ for each $\lambda^{*}$, the bias terms (in the lower panel for each $\lambda$) decrease substantially from the ones before, and they also decrease to stable values as we increase the number of iterations in the second-stage bias correction. Notably, the bias terms are all significantly smaller than those in the upper panel without the ridge-screening procedure. This is in agreement with our theoretical results in Theorem 5 and Theorem 7.
Table 5: Empirical averaging estimation errors (AEEs) when $p>n$ in Example 2, where the AEEs are similarly defined as those in (41) and (42) for ${\mathbf{b}}_{\lambda}$, ${\mathbf{b}}_{\lambda,k}$, and ${\mathbf{b}}_{\lambda,k,l}$. For each $\lambda$, the upper panel reports the AEES before ridge-screening, and the lower one presents the AEEs after ridge-screening with $k=100$ and $n^{*}=40$ in the variable selection. The number of iterations is set to $k=1,5,10,20,50$, and $100$ for the first-stage de-biased estimation, and $l=1,5,10,20,50,$ and $100$ for the second-stage bias-correction following the ridge-screening. 1000 replications are used in the experiments.
@@
We plot the empirical histograms of $\sqrt{n}\boldsymbol{\gamma}_{1}^{\prime}(\widehat{\boldsymbol{\beta}}(\lambda)-\boldsymbol{\beta})$, $\sqrt{n}\boldsymbol{\gamma}_{2}^{\prime}(\widehat{\boldsymbol{\beta}}(\lambda)-\boldsymbol{\beta})$, $\sqrt{n}\boldsymbol{\gamma}_{3}^{\prime}(\widehat{\boldsymbol{\beta}}(\lambda)-\boldsymbol{\beta})$, and $\sqrt{n}\boldsymbol{\gamma}_{4}^{\prime}(\widehat{\boldsymbol{\beta}}(\lambda)-\boldsymbol{\beta})$ in Figure 5, from which it is observed that most of the estimators are not centered around zero. Subsequently, we apply the bias-correction procedure to the original ridge estimators, and the empirical histograms of the bias-corrected estimators are presented in Figure 6 as compared to Figure 5. From Figure 6, it can be seen that while some estimators become more centered around zero (e.g., (a) and (b)) due to the bias-correction procedure, the validity of inference without ridge-screening cannot be assured, as (c) and (d) still exhibit significant biases.
Furthermore, we apply the ridge-screening approach to the de-biased estimators depicted in Figure 6, subsequently obtaining restricted ridge estimators along with their de-biased counterparts. Figure 7 showcases the empirical histograms of: (a) $\sqrt{n}\boldsymbol{\gamma}_{1}^{\prime}(\widehat{\boldsymbol{\beta}}_{\mathcal{M}_{k},l}^{*}(\lambda)-\boldsymbol{\beta})$; (b) $\sqrt{n}\boldsymbol{\gamma}_{2}^{\prime}(\widehat{\boldsymbol{\beta}}_{\mathcal{M}_{k},l}^{*}(\lambda)-\boldsymbol{\beta})$; (c) $\sqrt{n}\boldsymbol{\gamma}_{3}^{\prime}(\widehat{\boldsymbol{\beta}}_{\mathcal{M}_{k},l}^{*}(\lambda)-\boldsymbol{\beta})$; and (d) $\sqrt{n}\boldsymbol{\gamma}_{4}^{\prime}(\widehat{\boldsymbol{\beta}}_{\mathcal{M}_{k},l}^{*}(\lambda)-\boldsymbol{\beta})$, where $\widehat{\boldsymbol{\beta}}_{\mathcal{M}_{k},l}^{*}(\lambda)$ is a $p$-dimensional vector with $\widehat{\boldsymbol{\beta}}_{\mathcal{M}_{k},l}(\lambda)$ as its sub-vector and the remaining elements are set to zero. Similar to the curves in Figure 3, the density curve in Figure 7 is plotted based on the limiting distribution specified in Theorem 6(ii). From Figure 7, it is evident that the de-biased restricted ridge estimators predominantly cluster around zero. The density curves for the corresponding estimators are quite satisfactory, suggesting the validity of the proposed inference approach.
@@
4   Empirical Application
In this section, we apply the proposed method to forecast the U.S. macroeconomic series and study the out-of-sample prediction intervals using ridge regression. We consider the widely used macroeconomic variables studied by Stock and Watson (2002a), McCracken and Ng (2016), Giannone et al. (2021), and Gao and Tsay (2022, 2023), among many others. The data are obtained from the FRED-MD data base which are maintained by St. Louis Fed. See https://research.stlouisfed.org/econ/mccracken/fred-databases/. There are 127 variables in the online data set, and we remove 4 of them because of missing values therein. Consequently, we consider the 123 macro variables spanning from July 1962 to December 2019 as all the series have no missing values during this period, which is the same as the setting in Gao and Tsay (2023). The detailed variables and transformation codes to ensure the stationarity of each macro variable are provided in Table IA.I of Gao and Tsay (2023). The sample size is $n=690$.
Similar to Stock and Watson (2002b), Bai and Ng (2006), and Cheng and Hansen (2015), we adopt the following factor-augmented regression model:

### v2

@@ 3   Monte Carlo Simulations
We conduct Monte-Carlo experiments to demonstrate the performance of the proposed procedure by considering two scenarios where $p<n$ and $p>n$. Our goal is to study the effect of the bias-correction procedure and the approximation of the asymptotic normalities established in Section 2.5.
Example 1. In this example, we employ the data-generating process in (2) for different settings of $(p,n)$ with $p<n$. For each configuration of $(p,n)$, we set the seed number in R software to 1234 and generate an $n\times p$ matrix ${\mathbf{H}}$, where its elements are drawn independently from a Uniform distribution $U(-2,2)$. We then perform a singular-value decomposition on ${\mathbf{H}}$ and obtain its left and right singular vectors ${\mathbf{M}}\in R^{n\times p}$ and ${\mathbf{N}}\in R^{p\times p}$, then we let ${\mathbf{X}}={\mathbf{M}}{\mathbf{N}}^{\prime}$. The first $p/2$ elements of $\boldsymbol{\beta}$ are generated from $U(-2,-1)$ independently, and the remaining $p/2$ elements are from $U(1,2)$. In each replication, the noise $\varepsilon$ is generated from multivariate normal distribution $N({\bf 0},{\mathbf{I}}_{n})$. We consider the scenarios where $(p,n)=(50,100)$, $(50,400)$, $(100,200)$, and $(100,500)$. The choices of $\lambda$ are $\lambda=0.05n$, $0.1n$, $0.3n$, and $0.5n$. 1000 replications are used in each setting throughout the experiments.
We report the estimation errors of the ridge estimators and the de-biased ones in Table 1. For each configuration of $(p,n,\lambda)$, we define the empirical mean squared errors (MSEs) as
@@
where $\widehat{\boldsymbol{\beta}}^{(j)}(\lambda)$ and $\widehat{\boldsymbol{\beta}}_{c,k}^{(j)}(\lambda)$ are the ridge estimators and the de-biased one with $k$ iterations in the $j$-th replication, respectively. The standard errors $\widehat{\sigma}_{0}$ and $\widehat{\sigma}_{k}$ are estimated via
$\widehat{\sigma}_{0}=\sqrt{\frac{1}{1000}\sum_{j=1}^{1000}\frac{1}{n}\|{\mathbf{y}}-{\mathbf{X}}\widehat{\boldsymbol{\beta}}^{(j)}(\lambda)\|_{2}^{2}}\,\,\text{and}\,\,\widehat{\sigma}_{k}=\sqrt{\frac{1}{1000}\sum_{j=1}^{1000}\frac{1}{n}\|{\mathbf{y}}-{\mathbf{X}}\widehat{\boldsymbol{\beta}}_{c,k}^{(j)}(\lambda)\|_{2}^{2}}.$ (43)
Thus, we evaluate the estimations of the standard errors by the ridge estimators and the de-biased ones.
For each setting in Table 1, we see that the MSE first decreases as the number of iterations increases, but it increases if we use too many iterations to correct the bias terms. This is understandable because it is in line with the bias-variance trade-off in the machine learning literature. See, for example, Hastie et al. (2009). Specifically, we plot the MSEs for $(p,n)=(50,100)$ and $(100,200)$ with $\lambda=0.05n$ in Figure 1, where we can clearly see that the MSEs can achieve a minimum point as we increase the number of iterations, or equivalently, the model complexity, and the MSE will increase if we try to completely correct the biases. Finally, the MSE becomes stable, which is consistent with our asymptotic theory. Furthermore, comparing $\widehat{\sigma}$ and $\widehat{\sigma}_{100}$, we see that the standard error estimated using the debiased ridge estimator is closer to the true one (unity), showing the effectiveness of our bias-correction procedure.
@@
Next, we study the consistency of the de-biased estimators using the proposed method. For simplicity, we investigate the average estimation errors (AEEs) of the ridge and de-biased estimators. These AEE are defined, respectively, as
@@
based on the 1000 simulations. These measures, where we use $\ell_{2}$-norm to quantify the empirical biases. The AEEs are reported in Table 2. From the table, we can see a decreasing pattern for each $(p,n)$ as the number of iterations increases, implying that our de-biased estimators are consistent to the true parameters for sufficiently large $k$.
Table 2: Empirical average estimation errors (AEEs) when $p<n$ in Example 1, where $p$ is the number of predictors, $n$ is the sample size, and AEEs are defined in (44) and (45) for ${\mathbf{b}}_{\lambda}$ and ${\mathbf{b}}_{\lambda,k}$, respectively. The numbers of bias-correction iterations used are $k=1,5,10,20,50$, and $100$. The results are based on 1000 replications of the experiments.
@@
Finally, we investigate the performance of statistical inference using the de-biased estimators given in Theorem 6 for $p<n$. For simplicity, we only consider the case when $(p,n)=(100,200)$ and $\lambda=0.3n$, and can produce similar results for other settings. Let ${\mathbf{e}}_{i}$ be the $i$-th standard unit vector of $R^{p}$,
@@
where ${\bf 0}_{s}$ is an $s$-dimensional vector of zeros. Figure 2 presents the histograms of $\sqrt{n}{\mathbf{e}}_{1}^{\prime}(\widehat{\boldsymbol{\beta}}(\lambda)-\boldsymbol{\beta})$, $\sqrt{n}{\mathbf{e}}_{2}^{\prime}(\widehat{\boldsymbol{\beta}}(\lambda)-\boldsymbol{\beta})$, $\sqrt{n}\boldsymbol{\theta}_{1}^{\prime}(\widehat{\boldsymbol{\beta}}(\lambda)-\boldsymbol{\beta})$, and $\sqrt{n}\boldsymbol{\theta}_{2}^{\prime}(\widehat{\boldsymbol{\beta}}(\lambda)-\boldsymbol{\beta})$ obtained from 1000 experiments. From Figure 2, we see that the biases of the empirical means in all the histogram plots are significantly large and diverge from zero, implying that the traditional ridge estimators are not appropriate for making statistical inference as they overlook the bias terms. In contrast, we plot the empirical histograms of the de-biased estimators in Figure 3 under the same setting. From Figure 3, we see that the finite sample performance is quite satisfactory, and the bias effect of the de-biased estimators is small. In addition, the curve of the normal distribution in Theorem 6(i) is added to the corresponding histograms, where the standard error term $\sigma$ is estimated from the data as that in Remark 5(iv) and Table 1. From these curves, we can further confirm that the proposed inference method is valid for the de-biased ridge estimators, indicating that one can make valid inferences using the ridge estimators with our bias-correction method in practice.
@@
Figure 3: Empirical histograms of (a) $\sqrt{n}{\mathbf{e}}_{1}^{\prime}(\widehat{\boldsymbol{\beta}}_{c,k}(\lambda)-\boldsymbol{\beta})$; (b) $\sqrt{n}{\mathbf{e}}_{2}^{\prime}(\widehat{\boldsymbol{\beta}}_{c,k}(\lambda)-\boldsymbol{\beta})$; (c) $\sqrt{n}\boldsymbol{\theta}_{1}^{\prime}(\widehat{\boldsymbol{\beta}}_{c,k}(\lambda)-\boldsymbol{\beta})$; and (d) $\sqrt{n}\boldsymbol{\theta}_{2}^{\prime}(\widehat{\boldsymbol{\beta}}_{c,k}(\lambda)-\boldsymbol{\beta})$ in Example 1. We choose $k=120$ iterations for the de-biased estimators with $(p,n)=(100,200)$ and $\lambda=0.3n$. A normal density curve is plotted based on the limiting distribution outlined in Theorem 6(i). 1000 replications are used in the experiments.
@@
Example 2. In this example, we consider the scenario of $p>n$. The seed number used in this example is the same as that in Example 1. Each row of the design matrix ${\mathbf{X}}$ is generated independently from multivariate standard normal distribution $N({\bf 0},{\mathbf{I}}_{p})$. For each configuration of $(p,n)$, the first $5$ elements in $\boldsymbol{\beta}$ are generated from $U(-5,-2)$, the $6$th to the $10$th elements are from $U(2,5)$, and the remaining ones are zeros. In other words, we consider a sparse vector $\boldsymbol{\beta}$ with $p^{*}=10$ in (18). The dimensions and the sample sizes are $(p,n)=(150,120),(150,140),(220,180)$, and $(220,200)$, and we consider $\lambda=0.1n$, $0.3n$, and $0.8n$ in each replication for any given $(p,n)$. For simplicity, we choose $n^{*}=40$ in the ridge screening of (20) with $k=100$ iterations. $1000$ replications are used for each configuration of $(p,n)$ throughout the experiments.
@@
In Table 3, we report the empirical MSEs of the ridge estimators, the de-biased estimators before applying the ridge-screening method, and those after the ridge-screening with $k=100$ iterations in the first stage of bias-corrections. Specifically, the upper panel in Table 3 for each $\lambda$ presents the MSEs of the ridge estimators and the de-biased estimators with 100 iterations before applying the ridge-screening approach. We see that the MSEs of the de-biased estimators become quite stable as we increase the number of iterations. This result is consistent with our findings that there exists some minor bias that cannot be corrected from the data. In addition, the estimated standard errors are not close to unity, which is the true value in the experiments. In the lower panel of Table 3 for each $\lambda$, we apply the ridge-screening approach with $k=100$ iterations and sort out the largest $n^{*}=40$ components of $\widehat{\boldsymbol{\beta}}_{c,100}(\lambda)$ to obtain the restricted ridge estimators. We further apply the bias-correction approach, and the de-biased estimators are obtained in another $100$ iterations. From the lower panels of Table 3 for each $\lambda$, we see that the bias of the de-biased estimators decreases sharply to a relatively stable value for each configuration of $(p,n)$, which is understandable since the selected model consists of more parameters than those in the true one. In addition, the estimation errors are significantly smaller than those without the ridge-screening approach, and the estimated standard errors after bias correction are closer to the true one than those of the methods without using the ridge-screening and the bias-correction. We also note that the estimated standard errors are still larger than the true one (unity) when $\lambda=0.8n$ after the ridge-screening, but they are much closer to the true ones than those produced by the original ridge estimators. In addition, the result can be optimised by choosing more appropriate tuning parameter $\lambda$ (e.g. $\lambda=0.1n$ or $0.3n$) such that the estimated standard errors are close to one. Furthermore, we also note that there is also a bias-variance trade-off in the MSEs of the de-biased estimators with or without the ridge-screening. Figure A.1 in Sec. B of the Supplement plots the the empirical MSEs of the de-biased ridge estimators for different number of iterations for $\lambda=0.5n$ and $(p,n)=(150,120)$, where Part (a) plots the MSEs of the de-baised ridge estimators before applying the ridge-screening method, and Part (b) provides the MSEs of the de-biased ridge estimators after the ridge-screening. From the plots, it is clear that there is an optimal number of iteration that minimises the MSE, which is in line with our asymptotic results in Theorem 7.
@@
We also study the performance of the proposed ridge-screening method in Table A.1 of the Supplement, where the empirical probability (EP) is calculated by
@@
where $\mathcal{M}_{k}^{i}(\lambda^{*})$ is the recovery in the $i$-th experiment and $|\mathcal{M}_{0}\cap\mathcal{M}_{k}^{i}(\lambda^{*})|$ is the cardinality of the set $\mathcal{M}_{0}\cap\mathcal{M}_{k}^{i}(\lambda^{*})$. From Table A.1, the proposed method provides satisfactory performance in variable selections. The comparison results with other variable selection approaches, such as the LASSO, are not provided in this experiment due to the empirical probability of correct recoveries by the proposed RS method being $100\%$ in all settings of Table A.1.
@@
Similarly to the experiments in Example 1, we present the empirical average estimation errors (AEEs) in Table 4. The settings of $(p,n)$, $\lambda^{*}$, $\lambda$, $k$, and $l$ are the same as those in Table 3. From Table 4, we see that the average estimation errors before ridge-screening are also quite stable as we increase the number of iterations, which is similar to the findings in Table 3. After we apply the ridge-screening approach to the de-biased estimators in the upper panel of Table 4 with $k=100$ for each $\lambda^{*}$, the bias terms (in the lower panel for each $\lambda$) decrease substantially from their previous values, and the biases also decrease to stable values as we increase the number of iterations in the second-stage bias correction. Notably, the bias terms are all significantly smaller than those in the upper panel without the ridge-screening procedure. This finding is in agreement with our theoretical results in Theorem 5 and Theorem 7.
Table 4: Empirical average estimation errors (AEEs) when $p>n$ in Example 2, where the AEEs are similarly defined as those in (41) and (42) for ${\mathbf{b}}_{\lambda}$, ${\mathbf{b}}_{\lambda,k}$, and ${\mathbf{b}}_{\lambda,k,l}$. For each $\lambda$, the upper panel reports the AEES before ridge-screening, and the lower one presents the AEEs after ridge-screening with $k=100$ and $n^{*}=40$ in the variable selection. The number of iterations is set to $k=1,5,10,20,50$, and $100$ for the first-stage de-biased estimation, and $l=1,5,10,20,50,$ and $100$ for the second-stage bias-correction following the ridge-screening. 1000 replications are used in the experiments.
@@
We plot the empirical histograms of $\sqrt{n}\boldsymbol{\gamma}_{1}^{\prime}(\widehat{\boldsymbol{\beta}}(\lambda)-\boldsymbol{\beta})$, $\sqrt{n}\boldsymbol{\gamma}_{2}^{\prime}(\widehat{\boldsymbol{\beta}}(\lambda)-\boldsymbol{\beta})$, $\sqrt{n}\boldsymbol{\gamma}_{3}^{\prime}(\widehat{\boldsymbol{\beta}}(\lambda)-\boldsymbol{\beta})$, and $\sqrt{n}\boldsymbol{\gamma}_{4}^{\prime}(\widehat{\boldsymbol{\beta}}(\lambda)-\boldsymbol{\beta})$ in Figure 4, from which it is observed that most of the estimators are not centred around zero. Subsequently, we apply the bias-correction procedure to the original ridge estimators, and the empirical histograms of the bias-corrected estimators are presented in Figure 5 as compared to Figure 4. From Figure 5, it can be seen that while some estimators become more centred around zero (e.g., (a) and (b)) due to the bias-correction procedure, the validity of inference without ridge-screening cannot be assured, as (c) and (d) still exhibit significant biases.
Furthermore, we apply the ridge-screening approach to the de-biased estimators depicted in Figure 5, and subsequently obtain restricted ridge estimators along with their de-biased counterparts. Figure 6 showcases the empirical histograms of: (a) $\sqrt{n}\boldsymbol{\gamma}_{1}^{\prime}(\widehat{\boldsymbol{\beta}}_{\mathcal{M}_{k},l}^{*}(\lambda)-\boldsymbol{\beta})$; (b) $\sqrt{n}\boldsymbol{\gamma}_{2}^{\prime}(\widehat{\boldsymbol{\beta}}_{\mathcal{M}_{k},l}^{*}(\lambda)-\boldsymbol{\beta})$; (c) $\sqrt{n}\boldsymbol{\gamma}_{3}^{\prime}(\widehat{\boldsymbol{\beta}}_{\mathcal{M}_{k},l}^{*}(\lambda)-\boldsymbol{\beta})$; and (d) $\sqrt{n}\boldsymbol{\gamma}_{4}^{\prime}(\widehat{\boldsymbol{\beta}}_{\mathcal{M}_{k},l}^{*}(\lambda)-\boldsymbol{\beta})$, where $\widehat{\boldsymbol{\beta}}_{\mathcal{M}_{k},l}^{*}(\lambda)$ is a $p$-dimensional vector with $\widehat{\boldsymbol{\beta}}_{\mathcal{M}_{k},l}(\lambda)$ as its sub-vector and the remaining elements are zero. Similar to the curves in Figure 3, the density curve in Figure 6 is plotted based on the limiting distribution specified in Theorem 6(ii). From Figure 6, it is evident that the de-biased restricted ridge estimators predominantly cluster around zero. The density curves for the corresponding estimators are quite satisfactory, suggesting the validity of the proposed inference approach.
@@
4   An Empirical Application
In this section, we apply the proposed method to forecast the U.S. macroeconomic series and study the out-of-sample prediction intervals using ridge regression. We consider the widely used macroeconomic variables studied by Stock and Watson (2002a), McCracken and Ng (2016), Giannone et al. (2021), and Gao and Tsay (2022, 2024), among many others. The data are obtained from the FRED-MD database which is maintained by St. Louis Federal Reserve Bank. See https://research.stlouisfed.org/econ/mccracken/fred-databases/. There are 127 variables in the online data set, and we remove 4 of them because of missing values therein. Consequently, we consider the 123 macro variables spanning from July 1962 to December 2019 as all the series have no missing values during this period. The resulting data set is the same as that in Gao and Tsay (2024). The detailed variables and transformation codes to ensure the stationarity of each macro variable are provided in Table IA.I of Gao and Tsay (2024). The sample size is $n=690$.
Similarly to Stock and Watson (2002b), Bai and Ng (2006), and Cheng and Hansen (2015), we adopt the following factor-augmented regression model: